Program Changes as of Monday, March 25, 1:00 PM

Additions:

• Saturday 3:00 PM in 131: A. Ruzmaikina: Quasi Stationary States of the Indy 500 model

Title Changes:

• Tuesday 1:30 PM in 146: D. Gilbert: An inverse spectral problem for the one-dimensional Schrödinger operator

Cancellations:

• R. Brown (Tuesday 2:00 PM in 131)

• M. Juras (Wednesday 5:00 PM in 131)

• V. Kostrykin (Wednesday 5:00 PM in 146)

• H. Liu (Saturday 2:00 PM in 133)

• M. Tidiri (Saturday 3:00 PM in 133)

• I. Tralle (Tuesday 2:30 PM in 146)

• E. Tsekanovskii (Saturday 2:30 PM in 146)